Two electric train A and B leave the same station on parallel rail line train A start at rest with a uniform acceleration of 0.2m/s2 and attain a speed of 45km/h and maintain this speed at constant rate after ward . the train B leave the station 1 minute after train A has left and move with uniform acceleration of 0.4m/s2 to attain a maximum speed of 72 km/h. If train B maintain the speed of 72km/h determine the time it will over take train A

1
Expert's answer
2022-05-20T04:53:39-0400

"s_1=s_2,"

"v_1t+\frac{a_1t^2}2=v_2(t-1)+\frac{a_2(t-1)^2}2,"

"t=24~s."